A simple pendulum is a massless string with a point mass at the end. The mass is 0.6 kg. The length of the pendulum is 0.1 m. The amplitude is 7.5°. What is the period of the pendulum?
Answer #1

T=2*Pi*(L/g)1/2

L=0.1m

T=2*3.14*(0.1/9.8)1/2=0.634 seconds